OpenStudy (anonymous):

helpp please ill award you 3x + 14 = -2x x =

OpenStudy (anonymous):

-2x-3x=14 ==>-5x=14 ==> x=-14/5

OpenStudy (anonymous):

2(3 – y) = 6 – 2(1 – 3y) y = Last one of my test here

OpenStudy (anonymous):

your last one is y= 1/4